# Candidate Monotonicity and Proportionality for Lotteries and Non-Resolute Rules

### v1

@@ 4 A Probabilistic Version of Phragmén’s Ordered Rule
We need to show two things, (i) it respects the probabilities and (ii) the committees output by the rounding scheme satisfy PSC.
For the first point, we see that the probability spent on a single candidate $c$ is at most $\sum_{v\in V}p_{c}(v)$ by definition. Further, since for $r=m$ all voters form a solid coalition, it is also exactly $\sum_{v\in V}p_{c}(v)$ since $\sum_{c^{\prime}\in C,v\in V}p_{c}(v)$ is integral and only integral quantities get deducted. Thus, at the final step, there is still some integral probability mass left leading to all probabilities being fulfilled. Since each candidate has a probability mass of at most one, no candidate is elected more than once.
For (ii) let $N^{\prime}$ be a solid coalition over $C^{\prime}$ such that $\lvert N^{\prime}\rvert\geq\frac{\ell n}{k}$ for some $\ell\in[k]$. For $r=\lvert C^{\prime}\rvert$ we know that $C^{\prime}$ forms a prefix of the preferences of $N^{\prime}$. Any candidate from $C^{\prime}$ who has a probability of $1$ of getting selected is already included in the committee. Hence, the only remaining candidates have a probability of less than $1$. Let $C^{\prime\prime}$ be these candidates. By the way PFR runs, we know that the voters do not eat any candidate ranked worse. Hence, since the $C^{\prime}$ form a prefix for all voters in $N^{\prime}$ and since at most $\lvert C^{\prime}\setminus C^{\prime\prime}\rvert$ budget could have been spent on candidates with probability $1$, there is still a budget of at least $\frac{\lvert N^{\prime}\rvert k}{n}-\lvert C^{\prime}\setminus C^{\prime\prime}\rvert\geq\ell-\lvert C^{\prime}\setminus C^{\prime\prime}\rvert$ left at the start of the process. In iteration $r$ this will lead to the creation of vertices on the left side of the bipartition with weight at least $\lfloor\frac{\lvert N^{\prime}\rvert k}{n}-\lvert C^{\prime}\setminus C^{\prime\prime}\rvert\rfloor\geq\ell-\lvert C^{\prime}\setminus C^{\prime\prime}\rvert$. Since each vertex has integral weight, this means that either every candidate in $C^{\prime\prime}$ or at least $\ell-\lvert C^{\prime}\setminus C^{\prime\prime}\rvert$ candidates are selected from $C^{\prime}$ via dependent rounding, leading to at least $\min(C^{\prime},\ell)$ candidates being selected, therefore showing that PSC is satisfied. ∎
Adding on to this, we can however also show that PFR cannot be used to obtain the stronger guarantee of rank-PJR (Brill and Peters, 2023).
@@
Consider Example 5 of Brill and Peters (2023) with $4$ voters, $6$ candidates, and $k=2$:
@@
Here PFR would place a probability mass of $\frac{1}{2}$ on each of $c_{1},c_{4},c_{5},c_{6}$. This, however, is not compatible with rank-PJR, as it requires at least one of $c_{2}$ or $c_{3}$ to be selected.
@@ 4.1 Consequences for Party List Elections
A special case of multiwinner voting are so-called party list elections (Brill et al., 2024) in which a candidate can not only be elected once but multiple times. Of special interest is the apportionment problem. In it, voters give their vote for a single party, which then is aggregated into seats for the parties. That is, there are $m$ parties, with party $i$ receiving a $p_{i}$ share of the votes and $h$ seats to be distributed among the parties. Just as for multiwinner voting, there are also “proportionality” axioms for apportionment. Most prominently, the quota axiom states that party $i$ should receive either $\lfloor p_{i}h\rfloor$ or $\lceil p_{i}h\rceil$ seats, i.e., the number of seats for each party should be correctly rounded. Since allocation rules satisfying quota and other prominent desirable axioms do not exist (Balinski and Young, 2001), recent work has studied the probabilistic allocation of seats (Grimmet, 2004; Gölz et al., 2024; Correa et al., 2024): the vote shares of the parties get probabilistically rounded in such a way that each party receives their exact quota in expectation and always their rounded quota ex-post. Here, we shortly note that we can apply our result on PFR to show that if one additionally knew the complete rankings of each voter over all parties, one can get a randomized apportionment scheme satisfying the desirable quota properties in addition to PSC (when interpreting this apportionment instance as a multiwinner voting instance).
@@
For this, we note that PFR when run on a party-list instance, leads to probabilities exactly equal to the vote shares (as the candidates of a party never run out). Thus, Theorem 5 immediately gives us that ex-ante exact quota and PSC are satisfied. Further, due to the dependent rounding works, we also get that no party would be either over or underrepresented thus also giving us ex-post quota. ∎
@@
Figure 2: On the left, the number of instances with a given number of candidates for $k=3$ and $k=4$. On the right, the average number of candidates by instance for $k=4$ who have a $100\%$, at least $90\%$, at least $80\%$, at least $50\%$, or at most $10\%$ probability of getting elected under PFR.
@@
5 Experiments
We conduct some small experiments to measure how many committees our algorithm in Theorem 5 would actually randomize over. For this, we follow McCune and Graham-Squire (2024) and study their dataset of Scottish local government elections. This dataset consists of 1049 multiwinner voting instances with ranked preferences with on average 4800 voters, between 4 and 14 candidates and typically a $k$ of either $3$ or $4$. See the left side of Figure 2 for an overview over the distribution of $m$ for the different values of $k$. As the original preferences are not complete, we use the dataset of completed preferences constructed by Bardal et al. (2024). In their experimental work studying proportionality in this dataset, Bardal et al. (2024) already noted that in most instances, PSC is satisfied by most committees. In our experiments, we implement the dependent rounding scheme of Gandhi et al. (2006) and sample $50.000$ committees according to it for each instance. First, for the individual probabilities given by PFR, we see on the right side of  Figure 2 that typically on average one candidate is guaranteed to be selected. On the other hand, the number of candidates with a larger than $50\%$ probability is on average around $4$ for all values of $m$, while the average number of candidates with a very low chance (at most $10\%$) to be selected grows in the number of candidates, but even for $m=11$ does not exceed $4$, showing that usually there are several candidates with a non-trivial chance to be selected (the same graph for $k=3$ is very similar, see Figure 5 in the appendix).
Next, we measure the total number of committees with a positive probability and then the number of committees in the top $75\%$ (i.e., how many committees are needed to get a probability mass of at least $75\%$, if there were four committees sampled, each with probability $25\%$ this number would be $3$), $90\%$, $95\%$, and $99\%$ and compare these to the total number of committees satisfying PSC. We display these in Figure 4 for all values of $k$ and $m$ with more than $10$ instances. In it, we can see the effect already noticed by Bardal et al. (2024), most committees, especially for smaller values of $m$ satisfy PSC. Indeed, also most of these committees satisfying PSC were also sampled by us after fifty thousand samples. For the number of committees sampled with higher probabilities, one can for instance see that on average for $k=3$ and $m=7$ one of $7$ committees will be sampled with probability at least $75\%$ while on average with $99\%$ probability it will be one of $19$, with on average there being about $25$ out of the $35$ total committees satisfying PSC.
@@ 6 Conclusion
There are several open questions and directions for future work. Most importantly, whether there is a resolute voting rule satisfying PSC and candidate monotonicity is still open. Further, it is open whether the results presented here can be extended to stronger proportionality axioms (such as rank-PJR of Brill and Peters (2023)) or to the setting with weak preferences and the axiom of generalized PSC. For both, a characterization as provided by Aziz and Lee (2022) of the respective proportionality axioms is missing as well.
@@ 7 Acknowledgments
Appendix A Additional Graphic
[…]
Figure 5: The average number of candidates by instance for $k=3$ who have a $100\%$, at least $90\%$, at least $80\%$, at least $50\%$, or at most $10\%$ probability of getting elected under PFR.

### v2

@@ 4 A Probabilistic Version of Phragmén’s Ordered Rule
Besides this, we need to show two things, (i) this rounding scheme respects the probabilities and (ii) the committees output by the rounding scheme satisfy PSC.
For the first point, consider any candidate $c\in C$ and the vertices of $G$ connected to $c$. These vertices correspond to solid coalitions selected throughout the construction of $G$. By construction, for each voter in these solid coalitions, the probability that $c$ is selected is increased by at most $p_{c}(i)$. Hence, as the dependent rounding scheme respects the probabilities on the edges, $c$ is selected with a probability of at most $\sum_{i\in N}p_{c}(i)=p_{c}$.
Further, we know that for $r=m$ all voters form a solid coalition over all candidates. If at this step, there is any voter $i\in N$ left with $p_{c}(i)>0$, this probability will be part of the last vertex added to $G$. Thus, we know that $c$ is selected with a probability of exactly $p_{c}$.
[…]
Otherwise, let $C^{\prime\prime}$ be the set of candidates from $C^{\prime}$ with a probability of less than $1$ of being selected. As the voters in $N^{\prime}$ rank all candidates in $C^{\prime}$ before all candidates in $C\setminus C^{\prime}$ we know that no voter in $N^{\prime}$ can eat any candidate in $C\setminus C^{\prime}$ before all candidates in $C^{\prime}$ are assigned a probability of $1$. As we ran PFR until the total eaten weight was $k$, we get that the voters in $N^{\prime}$ consumed at least $\lvert N^{\prime}\rvert\frac{k}{n}\geq\ell$ many fractional candidates from $C^{\prime}$. For each rank in-between $1$ and $\lvert C^{\prime}\rvert$ voters in $N^{\prime}$ could have only taken part in solid coalitions containing exclusively candidates from $C^{\prime}$. Now consider $p^{\prime}\coloneqq\sum_{i\in N^{\prime}}\sum_{c\in C^{\prime}}p_{c}(i)$: the “budget” left at iteration $\lvert C^{\prime}\rvert$. As PFR only deducts integer values, we know that at the beginning of this iteration there must exist vertices in $G$ adjacent to a weight of at least $\ell-\lfloor p^{\prime}\rfloor$ in $G$ exclusively connected to candidates in $C^{\prime}$. In iteration $\lvert C^{\prime}\rvert$ PFR adds a vertex connected to candidates in $C^{\prime}$ with a total edge weight of $\lfloor p^{\prime}\rfloor$. As we round according to $G$ we know that each of these vertices will have a degree equal to the weight of their outgoing edges, with each of these vertices being included in the committee. Hence, we will select at least $\ell-\lfloor\sum_{i\in N^{\prime}}\sum_{c\in C^{\prime}}p_{c}(i)\rfloor+\lfloor\sum_{i\in N^{\prime}}\sum_{c\in C^{\prime}}p_{c}(i)\rfloor=\ell$ candidates from $C^{\prime}$, and therefore, the solid coalition $N^{\prime}$ is satisfies according to PSC. ∎
Adding on to this, we however also show that PFR cannot be used to obtain the recently introduced stronger guarantee of rank-PJR [BrPe23a]. To define rank-PJR, for any $i\in N$ and $c\in C$ we call $\operatorname{rank}(i,c)\coloneqq\lvert\{c^{\prime}\in C\colon c^{\prime}\succ_{i}c\}\rvert+1$ to be the rank given to $c$ by voter $i$ (i.e., voter $i$ gives candidate $c$ a rank of $r$ if this candidate appears in the $r$-th place of the voters rankings). Using this, we can define the property of rank-PJR.
Definition 6 (rank-PJR).
A committee $W$ satisfies rank-proportional justified representation (rank-PJR)[BrPe23a] if there is no group $N^{\prime}\subseteq N$ of voters, $C^{\prime}\subseteq C$ of candidates, $\ell\in[k]$, and rank $r\in[m]$ such that
  • i.)
    $\operatorname{rank}(i,c)\leq r$ for all $i\in N^{\prime}$ and $c\in C^{\prime}$
  • ii.)
    $\lvert N^{\prime}\rvert\geq\ell\frac{n}{k}$
  • iii.)
    $\lvert\{c\in W\colon\operatorname{rank}(i,c)\leq r\text{ for some }i\in N^{\prime}\}\rvert$$\ell$.
It is easy to see that any committee satisfying rank-PJR also satisfies PSC [BrPe23a]. Further, using an example of BrPe23a we can show that PFR cannot be decomposed to satisfy ex-post rank-PJR.
@@
Consider Example 5 of BrPe23a with $4$ voters, $6$ candidates, and $k=2$:
@@
Here, PFR would place a probability mass of $\frac{1}{2}$ on each of $c_{1},c_{4},c_{5},c_{6}$. This, however, is not compatible with rank-PJR. Independent of which two voters get their first choice candidate in the deterministic committee, the other two voters will rank candidate $c_{2}$ at a rank of at least $3$ while receiving no candidate at a rank of $3$, thus constituting a rank-PJR violation. In fact, as the PFR outcome is the only ex-ante PSC outcome in this instance, this also shows that ex-ante PSC and ex-post rank-PJR are incompatible.
@@ 4.1 Consequences for Party List Elections
A special case of multiwinner voting are so-called party list elections [BGP+24a] in which a candidate can not only be elected once but multiple times. Of special interest is the apportionment problem. In it, voters give their vote for a single party, which then is aggregated into seats for the parties. That is, there are $m$ parties, with party $i$ receiving a $p_{i}$ share of the votes and $h$ seats to be distributed among the parties. Just as for multiwinner voting, there are also “proportionality” axioms for apportionment. Most prominently, the quota axiom states that party $i$ should receive either $\lfloor p_{i}h\rfloor$ or $\lceil p_{i}h\rceil$ seats, i.e., the number of seats for each party should be correctly rounded. Since allocation rules satisfying quota and other prominent desirable axioms do not exist [BaYo01a], recent work has studied the probabilistic allocation of seats [Grim04a, GPP24a, JGU+24a]: the vote shares of the parties get probabilistically rounded in such a way that each party receives their exact quota in expectation and always their rounded quota ex-post. Here, we shortly note that we can apply our result on PFR to show that if one additionally knew the complete rankings of each voter over all parties, one can get a randomized apportionment scheme satisfying the desirable quota properties in addition to PSC (when interpreting this apportionment instance as a multiwinner voting instance).
@@
For this, we note that PFR when run on a party-list instance, leads to probabilities exactly equal to the vote shares (as the candidates of a party never run out). Thus, Theorem˜5 immediately gives us that ex-ante exact quota and PSC are satisfied. Further, due to the dependent rounding works, we also get that no party would be either over or underrepresented thus also giving us ex-post quota. ∎
44668810101212141405050100100150150Number of candidatesNumber of instancesk=3k=3k=4k=4
Figure 1: Number of instances with a given number of candidates for $k=3$ and $k=4$.
5 Experiments
We conduct some small experiments to measure how many committees our algorithm in Theorem˜5 would actually randomize over. For this, we follow McGr23a and study their dataset of Scottish local government elections. This dataset consists of 1049 multiwinner voting instances with ranked preferences with on average 4800 voters, between 4 and 14 candidates and typically a $k$ of either $3$ or $4$.
See Figure˜1 for an overview over the distribution of the number of instances for the different values of $k$ and $m$. As the original preferences are not complete, we use the dataset of completed preferences constructed by BBMP25a. In their experimental work studying proportionality in this dataset, BBMP25a already noted that in most instances, PSC is satisfied by almost all committees. In our experiments, we implement the dependent rounding scheme of GKPS06a and sample $50.000$ committees according to it for each instance. We are not aware of any computationally efficient way to extract these probabilities from the bipartite graph itself.
@@ 5 Experiments
Figure 3: Both plots depict the fraction of candidates who are assigned are assigned a probability in a certain range as a stackplot (for $k=3$ on the left and $k=4$ on the right).
First, for the individual probabilities given by PFR, we see in  Figure˜3 that typically — especially for larger values of $m$ — about $20\%$ to $40\%$ of the candidates have a probability of at least $70\%$ to get chosen by PFR. On the other hand, the ratio of candidates with a probability of less than $40\%$, is increasing with $m$ and even reaches $50\%$ for larger values of $m$. Thus, several candidates, typically have quite a low, yet non neglible, probability of getting elected.
Second, we measure the total number of committees with a positive probability and then the number of committees in the top $75\%$ (i.e., how many committees are needed to get a probability mass of at least $75\%$, if there were four committees sampled, each with probability $25\%$ this number would be $3$), $90\%$, $95\%$, and $99\%$ and compare these to the total number of committees satisfying PSC. We display these in Figure˜4 for all values of $k$ and $m$ with more than $10$ instances. In it, we can see the effect already noticed by BBMP25a, most committees, especially for smaller values of $m$ satisfy PSC. Indeed, also most of these committees satisfying PSC were also sampled by us after fifty thousand samples. In fact, as already observed by BBMP25a in several instances, barely any non-trivial solid coalitions exist. Thus, there are very few constraints on the output of PFR, which in turn leads to most committees being able to be selected. For the number of committees sampled with higher probabilities, one can for instance see that on average for $k=3$ and $m=7$ one of $7$ committees will be sampled with probability at least $75\%$ while on average with $99\%$ probability it will be one of $19$, with on average there being about $25$ out of the $35$ total committees satisfying PSC. We further note that the committee with the largest probability of being selected usually has a probability between $15\%$ and $35\%$.
@@
Thus, while PFR offers an appealing theoretical solution, it also inherits several key weaknesses of PSC as an axiom. Namely, due to its dependence on solid coalitions, the relative lack of meaningful solid coalitions [BBMP25a] in-turn leads to a lack of constraints for PFR, which still enables a large fraction of all PSC committees to be chosen by PFR. Hence, if one would want near-certainty on the outcome of the lottery, PFR is most likely not a good choice as an election rule. If, however, one is content with its use as a lottery, PFR might provide an interesting monotone alternative to classical deterministic voting rules.
@@ 6 Conclusion
There are several open questions and directions for future work. Most importantly, whether there is a resolute voting rule satisfying PSC and candidate monotonicity is still open. Further, it is open whether the results presented here can be extended to stronger proportionality axioms (such as rank-PJR of BrPe23a) or to the setting with weak preferences and the axiom of generalized PSC. For both, a characterization as provided by AzLe22a of the respective proportionality axioms is missing as well.
@@ 7 Acknowledgments
8 Compliance with Ethical Standards
This work was partially supported by the Singapore Ministry of Education under grant number MOE-T2EP20221-0001. No human participants or animals were used for this paper.